# Microlens-enhanced SiPMs

### v1

@@ Abstract
A novel concept to enhance the photo-detection efficiency (PDE) of silicon photomultipliers (SiPMs) has been applied and remarkable positive results can be reported. This concept uses arrays of microlenses to cover every second SiPM pixel in a checkerboard arrangement and aims to deflect the light from the dead region of the pixelised structure towards the active region in the center of the pixel. The PDE is improved up to 24%, external cross-talk is reduced by 40% compared to a flat epoxy layer, and single photon time resolution is improved. This detector development is conducted in the context of the next generation LHCb scintillating fibre tracker located in a high radiation environment with a total of 700’000 detector channels. The simulation and measurement results are in good agreement and will be discussed in this work.
@@ I Introduction
In the context of the scintillating fibre tracker (SciFi Tracker) [1], our group, the High Energy Physics Laboratory (LPHE) at EPFL, is working to improve the silicon photomultiplier (SiPM) photodetectors in order to overcome the challenges imposed by radiation, detector size, and the low material budget requirement for the next upgrade in 2034. The ultimate goal for this development is to improve the signal-to-noise ratio. We found that the enhancement of the SiPM with microlenses not only significantly increases the signal but also reduces external cross-talk and has the potential to reduce noise.
The choice of the ideal SiPM pixel size for a custom design is typically guided by the compromise between high geometrical fill factor (GFF) and the resulting high photo-detection efficiency (PDE) for large pixels and the manifold disadvantages resulting from higher gain and therefore higher correlated noise, longer recovery time, lower dynamic range, larger bias current. Adding the radiation environment as an additional criterion, small pixel size and low excess bias voltage (over-voltage, $\Delta V$) are also favoured in order to limit the effect of increasing dark count rate (DCR) and correlated noise. As a summary, small pixel size and low $\Delta V$ are advantageous in all aspects except for PDE.
The hit detection efficiency of the scintillating fibre tracker technology depends on the actual signal (scintillating light) produced and the PDE of the photodetector. At a given fibre technology and fibre length, stacking several layers of fibres is used to increase the signal for a traversing particle. This comes at the cost of higher material budget, large signal spread for incident particles with large angle and last but not least, higher cost for more scintillating fibres. The implementation of microlens arrays (MLAs) aligned with the pixel structure has been successfully tested and the results are presented in the following.
@@ II The microlens on SiPMs
Enhancing imaging photodetectors with pixel level microlenses is commonly used for CCD and CMOS camera sensors [7] and has also been successfully applied to SPAD array sensors in the last years [3, 5, 13]. A good overview of the fabrication process of refractive MLAs is given in [6], focusing on spherical lenses, although cylindrical lenses can be obtained with the same process [4]. The process used for our application is based on photoresist processing, followed by a thermal reflow yielding a master mold. This mold is used with a UV-curable polymer to replicate the MLA on the wafer level. Other fabrication processes are microdroplet inkjet printing or MEMS-based. The microlens implementation requires high-precision manufacturing, alignment and a clean-room environment. These types of processes are typically performed by highly specialised research institutions. For this project the Swiss Center for Electronics and Microtechnology (CSEM)  [^1] https://www.csem.ch was subcontracted for the microlens implementation.
Unlike the typical implementation where the base surface of the lens is square or hexagonal and placed to cover the entire sensor surface, we found an innovative solution by covering every second pixel in a checkerboard structure as shown in Fig. 1. As the SiPM is operated as a photon counter and not as a pixelised imaging sensor, the deviation of photons from one pixel to a neighbouring pixel is acceptable.
@@
In this arrangement, the microlens deflects the light by refraction, from the dead border region towards the center of the lens covered pixel. The area between lenses is covered with a flat region aligned to the active region of adjacent pixels. As illustrated by the pixel layouts in Fig. 2, the active region is reduced by optical insulation trenches (to suppress optical cross-talk), quench resistors, and metal traces. With an optical microscope, the GFFs of different pixel layouts can be measured and used in the simulation. For the shown examples, the values correspond well with the specifications given by the manufacturers Hamamatsu (HPK)  [^2] https://www.hamamatsu.com and Fondazione Bruno Kessler (FBK)  [^3] https://sd.fbk.eu.
@@
The second important modification due to the microlens is related to the concentration of the light towards the center of the pixels. At low $\Delta V$ and therefore low electric field, the region close to the pixel border has a low avalanche trigger probability ($P_{AT}$) and therefore a low PDE. With the microlens, the boarder region of the pixel is less used (light deviated towards the center) and as a consequence, PDE is enhanced at low $\Delta V$ [8]. Furthermore, the time resolution is also improved. In [2] it is shown that the single photon time resolution (SPTR) is improved for the center region of the pixel. Masking pixel boundaries has been successfully performed by FBK to improve SPTR [9]. With the effect of the lens, improved SPTR for the microlens-enhanced detectors is observed as for the masked structures.
@@ III Simulation
To optimise the geometry and to assess the performance of the microlens implementation, a custom simulation based on a ray tracing Monte Carlo using the ROOT framework [15] has been developed. The calculation of the effectiveness of the microlens is evaluated for a system of nine pixels with one microlens placed on top of the central pixel. For the actual implementation not only one in nine but every second pixel is covered with a lens and therefore the improvement expected from the implementation is scaled with a factor of 4.5 with respect to the simulation. The factor 4.5 corresponds to the mean number of lenses that can be implemented on a $3\times 3$ pixel grid. The simulation results were confirmed in an early phase with a commercial ray tracing simulation software [16] by our industrial partner CSEM. The functionality of the pixel is modelled as active area (center), dead border regions (outermost borders), and partially efficient low field border region (between dead and active). The GFF is set to a constant value in the simulation but experimental data show lower avalanche trigger probability at the edges [17]. Therefore, the GFF of the pixel was used to define the dead region size and the low field region was modeled according to the best estimation by the designer at FBK as a band with reduced PDE. The improvement is defined as an increase in detected photons. It is equal to the ratio between detected photons with a MLA and a flat coated SiPM devices. An illustration of the total incident photons, the detected photons, and the lost photons is given in Fig. 3.
Figure 3: (a) Illustration of the parameters of the lens and the detection regions of the SiPM pixel. (b) Distribution of the photons arriving at the silicon surface. Note that only a single lens in the center is present in the simulation. (c) Detected photon distribution. (d) Lost photons falling in the dead or partially efficient region. The dead regions with the microlens have a significant lower number of lost photons compared to the dead regions without microlens. The light source angular distribution used for (b), (c) and (d) corresponds to a $230~{}cm$ long SciFi.
@@ III-A Light source exit angle distribution
For the simulation, two types of light sources were used. One modelling the SciFi exit angle distribution and one with normal incident light as used in our detector characterisation setup. To obtain the exit angle distribution from the SciFi, a dedicated setup was used. A 2.5 m long fibre is excited with a UV-LED at a distance of 20 and 230cm from the fibre end. With a camera fixed on a rotary motorised stage, the intensity as a function of angle was measured and normalised. A $3~{}mm$ aperture placed at a distance of $5~{}cm$ from the fibre end ensures a small angular acceptance. Note that the distribution depends on the injection distance and the quality of the exit surface. In Fig. 4 the distribution for a SciFi is given. The long transmission path (long for a plastic optical fibre) reveals a higher attenuation for large angles.
@@
In the context of the improvement of the SciFi tracking application, the optimisation of the microlens parameters was performed with the SciFi exit angle distribution. SiPMs used in other applications such as for Cherenkov detectors [10], astroparticle detectors [11], and Lidar [12] require microlens parameters optimised to the specific light sources. In general, these aforementioned applications have narrow normal incident light sources.
@@ III-B The optical parameters
The simulation allowed to obtain the optimal geometrical parameters for the microlens. The range of the parameters was guided (limited) by the manufacturing process limitation for the microlens. The maximal value of the lens diameter is the full diagonal of a pixel but in the fabrication process of the mold, a minimal distance of $2\,\mu m$ between lenses is required. The simulation showed that the highest light detection is achieved with a value of 95% of the pixel diagonal (Table I). As this value is expressed as a percent of the pixel diagonal, it is valid for different pixel sizes.
The 95% results in a lens spacing of $2.95\,\mu m$ for a pixel size of $42\,\mu m$ and $2.4\,\mu m$ for a pixel size of $31\,\mu m$. The lens surface curvature is restricted to a spherical shape due to the manufacturing process of the master mold using the surface tension during a melting cycle. The height of the lens is restricted via the maximal acceptable inclination angle of the surface. The de-moulding process can only be reliably performed for angles below $75^{\circ}$ and therefore this limits the maximal lens height. For the simulation this angle is used for the parametrisation of the lens height. The third restriction applies to residual layer height ($RLT$), the thickness of the coating without lenses as illustrated in Fig. 3. With a large mold size covering an entire silicon wafer, the minimal $RTL$ that can be reached with this replication process is of the order of 10 to $15\,\mu m$.
@@ III-C Simulation results
In the simulation, $R_{Lens}$, $H_{Lens}$ and $RLT$ were varied to achieve maximum light yield enhancement ($LY_{enh}$). The results are summarized in Table I. For this study, the dead region of the pixel was uniformly distributed along the edge of the pixel. In addition, an intermediate region $2\,\mu m$ wide at the edge of the active region is implemented modeling the low field region (LFR) which is more pronounced at low $\Delta V$. Here, the efficiency is decreased from $100\%$ in Table I to $60\%$ in Table II where only $RLT$ is vaired.
@@ III-D Simulation beyond the current detector implementation and for normal incident light
In addition to the $41.7\,\mu m$ pixel from FBK with a very high $GFF=81.5\%$ and the large exit angle distribution light source (SciFi), other cases were studied. Particularly interesting is the optimisation of a narrow and normal incident light source. A list with parameters and simulated $LY_{enh}$ is given in Table III. There, a comparison of different pixels shown in Fig. 2 is given. The larger (SciFi) exit angle distribution leads to optimal parameters with small $RLT$ and high lenses ($H_{Lens}=75^{\circ}$). With a narrow and normal incident light source thicker $RLT$ and flatter lenses can be efficient. Smaller pixels and in general pixels with lower $GFF$ have the potential for a larger improvement. To simplify the comparison we assume a $2\,\mu m$ low field region with $R_{Lens}=95\%$ in all cases.
@@ III-E Optical coupling
The microlens principle is based on light refraction between air and the lens material. Any application that requires optical glue or grease between the light source and the photodetector (e.g. LYSO crystals glued to SiPM surface) will not profit from the microlens, whereas systems where a narrow exit angle distribution of the light source is present have a high potential for enhancing the PDE with MLAs.
@@ IV The implementation of prototypes
In view of the development of a new multichannel SiPM array for the LHCb Upgrade 2, a dedicated implementation was defined to evaluate the microlens enhancement. In addition to the constraint of the geometrical aspects of the active surface and the chip dimensions, additional specific alignment markers were added. To reach thin $RLT$, the original 8-inch wafers were diced into pieces of $2\times 2$ reticles of $32\,mm\times 32\,mm$ in size. The detectors with $41.7\,\mu m$ pixel size were chosen for this study and two types of samples were prepared, flat layer (uniform epoxy layer) and MLAs.
@@ IV-A Silicon
A custom implementation of a SiPM array was produced by FBK based on the NUV-HD technology [14]. This technology features a high $GFF$ of $81.5\%$ for a pixel size of $41.7\,\mu m$ and can be operated at a large $\Delta V$ range. For this production, three different pixel sizes were included in a reticle, but for what concerns the MLA implementation, only the $41.7\mu m$ chip type was used. The silicon dies are multichannel arrays with 64-channels with a size of $1.62\,mm\times 250\,\mu m$ per channel and a total size of $16\,mm\times 1.62\,mm$. For a MLA replication, 4 dies were produced at a single replication process but as other dies were also present on the reticle, the total replication area on a $2\times 2$ reticle is $32\,mm\times 32\,mm$. The connectivity is implemented with a common cathode on the back and bond pads on the top at a distance of $700\,\mu m$ from the active surface to allow for a bond wire protection with glue. The MLA layer deposition is restricted to the active surface but leaves the bond pad surface uncoated (Fig. 5).
@@
The deposit of the MLAs and flat layers was performed on eight dies each. The samples were evaluated for their optical properties before dicing and packaging. The package is a bare die bonding of two dies on a flexible Kapton circuit with a ceramic stiffener and temperature sensor on the backside. The assembled module can be seen in Fig. 6.
@@
For the electrical tests, these samples were characterized with a tunable monochromatic light source for the test with narrow normal incident light as well as with a SciFi mat (array of fibres) that provides the specific angular distribution already discussed in Sec. III-A.
@@ V Visual evaluation
Before the electrical properties were measured a visual inspection of the characteristics of the different samples was performed. The visual inspection was performed at several steps in the production of the mold and at the replication stage.

### v2

@@ Abstract
A novel concept to enhance the photo-detection efficiency (PDE) of silicon photomultipliers (SiPMs) has been applied and remarkable positive results can be reported. This concept uses arrays of microlenses to cover every second SiPM pixel in a chequerboard arrangement and aims to deflect the light from the dead region of the pixelised structure towards the active region in the centre of the pixel. The PDE is improved up to 24%, external cross-talk is reduced by 40% compared to a flat epoxy layer, and single photon time resolution is improved. This detector development is conducted in the context of the next generation LHCb scintillating fibre tracker located in a high radiation environment with a total of 700’000 detector channels. The simulation and measurement results are in good agreement and will be discussed in this work.
@@ I Introduction
In the context of the scintillating fibre tracker (SciFi Tracker) [1], our group, the High Energy Physics Laboratory (LPHE) at EPFL, is working to improve the silicon photomultiplier (SiPM) photodetectors in order to overcome the challenges imposed by radiation, detector size, and the low material budget requirement for the next upgrade in 2034. The ultimate goal of this development is to improve the signal-to-noise ratio. We found that the enhancement of the SiPM with microlenses not only significantly increases the signal but also reduces external cross-talk and has the potential to reduce noise.
The choice of the ideal size of the SiPM pixel for a custom design is typically guided by the compromise between the high geometrical fill factor (GFF) and the resulting high photo-detection efficiency (PDE) for large pixels and the manifold disadvantages resulting from higher gain and therefore higher correlated noise, longer recovery time, lower dynamic range, and higher bias current. Adding the radiation environment as an additional criterion, small pixel size and low excess bias voltage (over-voltage, $\Delta V$) are also preferred to reduce the effect of increasing the dark count rate (DCR) (heat dissipation and self-heating), as well as reducing correlated noise. In summary, the small pixel size and the low $\Delta V$ are advantageous in all aspects except for the PDE.
The hit detection efficiency of the scintillating fibre tracker technology depends on the actual signal (scintillating light) produced and the PDE of the photodetector. For a given fibre technology and fibre length, stacking several layers of fibres is used to increase the signal for a traversing particle. This comes at the cost of a higher material budget, a large signal spread for incident particles with a large incident angle, and last but not least a higher cost for more scintillating fibres. The implementation of microlens arrays (MLAs) aligned with the pixel structure has been successfully tested, and the results are presented in the following.
@@ II The microlens on SiPMs
Enhancing imaging photodetectors with pixel level microlenses is commonly used for CCD and CMOS camera sensors [2] and has also been successfully applied to SPAD array sensors in the last years [3, 4, 5]. A good overview of the fabrication process of refractive MLAs is given in [6], focusing on spherical lenses, although cylindrical lenses can be obtained with the same process [7]. The process used for our application is based on photoresist processing, followed by a thermal reflow yielding a master mold. This mold is used with a UV-curable polymer to replicate the MLA on the wafer level. Other fabrication processes are microdroplet inkjet printing or MEMS-based. The implementation of microlenses requires high-precision manufacturing, alignment, and a clean-room environment. These types of processes are typically performed by highly specialised research institutions. For this project, the Swiss Center for Electronics and Microtechnology (CSEM) [8] was sub-contracted for the microlens implementation.
Unlike the typical implementation where the base surface of the lens is square or hexagonal and placed to cover the entire sensor surface, we found an innovative solution by covering every second pixel in a chequerboard structure, as shown in Fig. 1. As the SiPM is operated as a photon counter and not as a pixelised imaging sensor, the deviation of photons from one pixel to a neighbouring pixel is acceptable.
@@
In this arrangement, the microlens deflects the light by refraction, from the dead border region towards the centre of the lens covered pixel. The area between lenses is covered with a flat region aligned with the active region of adjacent pixels. As illustrated by the pixel layouts in Fig. 2, the active region is reduced by optical insulation trenches (to suppress optical cross-talk), quench resistors, and metal traces. With an optical microscope, the GFFs of different pixel layouts can be measured and used in the simulation. For the shown examples, the values correspond well with the specifications given by the manufacturers Hamamatsu (HPK) [9] and Fondazione Bruno Kessler (FBK) [10].
@@
The second important modification due to the microlens is related to the concentration of light towards the centre of the pixels. At low $\Delta V$ and therefore low electric field, the region close to the pixel border has a low avalanche trigger probability ($P_{AT}$) and therefore a low PDE. With the microlens, the border region of the pixel is less used (light deviated towards the center) and, as a consequence, the PDE is expected to be enhanced at low $\Delta V$. This is supported by the simulation performed in Ref. [11] and as well shown in [12] by measuring the GFF as a function of $\Delta V$. Furthermore, the time resolution is also improved. In [13] it is shown that the single photon time resolution (SPTR) is improved for the center region of the pixel. Masking pixel boundaries has been successfully performed by FBK to improve SPTR [14]. With the effect of the lens, improved SPTR for the microlens-enhanced detectors is observed as for the masked structures.
@@ III Simulation
To optimise the geometry and assess the performance of the microlens implementation, a custom simulation based on a ray tracing Monte Carlo using the ROOT framework [15] has been developed. The calculation of the effectiveness of the microlens is evaluated for a system of nine pixels with one microlens placed on top of the central pixel. For the actual implementation, not only one in nine but every second pixel is covered with a lens, and therefore the improvement expected from the implementation is scaled with a factor of 4.5 with respect to the simulation. The factor 4.5 corresponds to the mean number of lenses that can be implemented on a grid of $3\times 3$ pixels. The simulation results were confirmed in an early phase with a commercial ray tracing simulation software [16] by our industrial partner CSEM. The functionality of the pixel is modelled as active area (centre), dead border regions (outermost borders), and partially efficient low field border region (between dead and active). The GFF is set to a constant value in the simulation but experimental data show lower avalanche trigger probability at the edges [12]. Therefore, the GFF of the pixel was used to define the dead region size and the low field region was modelled according to the best estimation by the designer at FBK as a band with reduced PDE. The improvement is defined as an increase in the number of detected photons. It is equal to the ratio between detected photons with a MLA and a flat coated SiPM device. An illustration of the total incident photons, the detected photons, and the lost photons is given in Fig. 3.
Figure 3: (a) Illustration of the parameters of the lens and the detection regions of the SiPM pixel. (b) Distribution of the photons arriving at the silicon surface. Note that only a single lens in the center is present in the simulation. (c) Detected photon distribution. (d) Lost photons falling in the dead or partially efficient region. The dead regions with the microlens have a significant lower number of lost photons compared to the dead regions without microlens. The light source angular distribution used for (b), (c) and (d) corresponds to a $230~{}cm$ long SciFi (polished).
@@ III-A Light source exit angle distribution
For the simulation, two types of light sources were used. One modelling the SciFi exit angle distribution and one with normal incident light as used in our detector characterisation setup. To obtain the exit angle distribution from the SciFi, a dedicated setup was used. A 2.5 m long fibre is excited with a UV-LED at a distance of 20 and 230cm from the fibre end. With a camera fixed on a rotary motorised stage, the intensity as a function of angle was measured and normalised. A $3~{}mm$ aperture placed at a distance of $5~{}cm$ from the end of the fibre ensures a small angular acceptance. Note that the distribution depends on the injection distance and the quality of the exit surface. In Fig. 4 the distribution for a SciFi is given. The long transmission path (long for a plastic optical fibre) reveals a higher attenuation for large angles.
@@
In the context of the improvement of the SciFi tracking application, the optimisation of the microlens parameters was performed with the SciFi exit angle distribution. SiPMs used in other applications, such as for Cherenkov detectors [17], astroparticle detectors [18], and Lidar [19], require microlens parameters optimised for specific light sources. In general, these aforementioned applications have narrow normal incident light sources.
@@ III-B The optical parameters
The simulation allowed us to obtain the optimal geometrical parameters for the microlens. The range of parameters was guided (limited) by the limitations of the microlens manufacturing process. The maximal value of the diameter of the lens is the full diagonal of a pixel, but in the mold fabrication process, a minimal distance of $2\,\mu m$ between the lenses is required. The simulation showed that the highest light detection is achieved with a value of 95% of the pixel diagonal (Table I). As this value is expressed as a percentage of the pixel diagonal, it is valid for different pixel sizes.
The 95% results in a lens spacing of $2.95\,\mu m$ for a pixel size of $42\,\mu m$ and $2.4\,\mu m$ for a pixel size of $31\,\mu m$. The lens surface curvature is restricted to a spherical shape due to the manufacturing process of the master mould that uses the surface tension during a melting cycle. The height of the lens is restricted via the maximum acceptable inclination angle of the surface. The demolding process can only be reliably performed for angles below $75^{\circ}$ and therefore this limits the maximal lens height. For the simulation, this angle is used for the parameterisation of the lens height. The third restriction applies to residual layer height ($RLT$), the thickness of the coating without lenses, as illustrated in Fig. 3. With a large mould size covering an entire silicon wafer, the minimal $RTL$ that can be reached with this replication process is on the order of 10 to $15\,\mu m$.
@@ III-C Simulation results
In the simulation, $R_{Lens}$, $H_{Lens}$, and $RLT$ were varied to achieve the highest enhancement in light yield ($LY_{enh}$). The results are summarised in Table I. For this study, the dead region of the pixel was uniformly distributed along the edge of the pixel. In addition, an intermediate region $2\,\mu m$ wide at the edge of the active region is implemented modeling the low field region (LFR) which is more pronounced at low $\Delta V$. Here, the efficiency decreases from $100\%$ in Table I to $60\%$ in Table II where only $RLT$ is varied.
@@ III-D Simulation beyond the current detector implementation and for normal incident light
In addition to the $41.7\,\mu m$ pixel from FBK with a very high $GFF=81.5\%$ and the large exit angle distribution light source (SciFi), other cases were studied. Particularly interesting is the optimisation of a narrow and normal incident light source. A list of parameters and simulated $LY_{enh}$ is given in Table III. There, a comparison of different pixels shown in Fig. 2 is given. The larger (SciFi) exit angle distribution leads to optimal parameters with small $RLT$ and high lenses ($H_{Lens}=75^{\circ}$). With a narrow and normal incident light source, thicker $RLT$ and flatter lenses can be efficient. Smaller pixels and in general pixels with lower $GFF$ have the potential for a larger improvement. To simplify the comparison, we assume a $2\,\mu m$ low field region with $R_{Lens}=95\%$ in all cases.
@@ III-E Optical coupling
The microlens principle is based on the refraction of light between the air and the lens material. Any application that requires optical glue or grease between the light source and the photodetector (e.g. LYSO crystals glued to SiPM surface) will not profit from the microlens, whereas systems where a narrow exit angle distribution of the light source is present have high potential for enhancing the PDE with MLAs.
@@ IV The implementation of prototypes
In view of the development of a new multichannel SiPM array for LHCb Upgrade 2, a dedicated implementation was defined to evaluate microlens enhancement. In addition to the constraint of the geometrical aspects of the active surface and the chip dimensions, additional specific alignment markers were added. To reach thin $RLT$, the original 8-inch wafers were diced into pieces of $2\times 2$ reticles of $32\,mm\times 32\,mm$ in size. The detectors with $41.7\,\mu m$ pixel size were chosen for this study and two types of samples were prepared, flat layer (uniform epoxy layer) and MLA.
@@ IV-A Silicon
A custom implementation of a SiPM array was produced by FBK based on the NUV-HD technology [20]. This technology features a high $GFF$ of $81.5\%$ for a pixel size of $41.7\,\mu m$ and can be operated in a wide $\Delta V$ range. For this production, three different pixel sizes were included in a reticle, but for the MLA implementation, only the $41.7\mu m$ chip type was used. The silicon dies are multichannel arrays with 64 channels with a size of $1.62\,mm\times 250\,\mu m$ per channel and a total size of $16\,mm\times 1.62\,mm$. For an MLA replication, 4 dies were produced in a single replication process, but since other dies were also present in the reticle, the total replication area on a $2\times 2$ reticle is $32\,mm\times 32\,mm$. The connectivity is implemented with a common cathode on the back and bond pads on the top at a distance of $700\,\mu m$ from the active surface to allow for a bond wire protection with glue. The MLA layer deposition is restricted to the active surface but leaves the surface of the bond pad uncoated (Fig. 5).
@@
The deposit of the MLAs and flat layers was performed on eight dies each. The optical properties of the samples were evaluated before dicing and packaging. The package is a bare die bonding of two dies on a flexible Kapton circuit with a ceramic stiffener and temperature sensor on the backside. The assembled module can be seen in Fig. 6.
@@
For electrical tests, these samples were characterised with a tunable monochromatic light source for the test with narrow normal incident light as well as with a SciFi mat (array of fibres) that provides the specific angular distribution discussed in Sect. III-A.
@@ V Visual evaluation
Before the electrical properties were measured, a visual inspection of the characteristics of the different samples was performed. Visual inspection was performed in several steps in mould production and in the replication stage.